Find two numbers whose sum is 114 and difference is 58?
Answer by DrBeeee(684)

Let the two numbers be a and b then you have
(1) a + b = 118 and
(2) a - b = 58
Add (1) and (2) and get
(3) 2a = 172 or
(4) a = 86
Then from (1) we have
(5) 86 + b = 118 or
(6) b = 28
Check this with (2)
Is (86 - 28 = 58)?
Is (58 = 58)? Yes
Answer: the two numbers are 58 and 86.